OpenStudy (anonymous):

I need to combine like expressions and have no idea how. 1/3 log_a⁡x+1/3 log_a⁡y+log_a⁡y

OpenStudy (amistre64):

kinda hard to parse the equation

OpenStudy (anonymous):

\[\log (xy) = logx + logy \]\[\log {x \over y} = \log x - \log y\] does this help?